# A continuous symmetry breaking measure for finite clusters using Jensen-Shannon divergence

### v1

@@ 5.1.1 The symmetry breaking from the boundary of the finite cluster
  • Spheroidal Cutout: Centered on a central atom, extending two unit cells along the major axis and one unit cell along each minor axis, containing all atoms within this volume.
@@
In this test, the lattice parameter for nickel was chosen to be $3.52$. The atomic displacement parameters (ADP) for each nickel atom were set to $U_{\text{iso}}$ $=0.013$2. For each nickel atom, the electron count is 28, which aligns with its atomic number, $Z$. Additionally, the occupancy, $o$, is set to one for each site. For the evaluation of SBM, both $\mathscr{S}^{KL}_{R_{\alpha}}[\mu]$ and $\mathscr{S}^{JS}_{R_{\alpha}}[\mu]$ were estimated using Monte Carlo random sampling. The sample sizes for these estimations were determined based on a 95% confidence interval, with a bilateral error tolerance set at $0.025$ for the KL-SBM and $0.0025$ for the JS-SBM.
@@
We first consider the point-symmetry-preserving cutouts in Fig. 2 (a and c). By definition, both $\mathscr{S}_{R_{\alpha}}[\mu]=0$ when $\alpha=0$ as the object is compared with itself. As we rotate the angle away from zero, $\mathscr{S}_{R_{\alpha}}[\mu]$ increase. Because the object has 4-fold rotational symmetry, $C_{4}$, for rotations about this axis, we expect both SBM $\mathscr{S}_{R_{\alpha}}=\mathscr{S}_{C_{4}^{i}}=0$ when $\alpha=i\pi/2$ rad for $i=1,2,3$, which results in $\mathscr{S}_{R_{\alpha}}$ being a function with 4-fold periodicity in $\alpha$, as is seen. The fact that these cutouts do not break the point-symmetry is evident in the fact that $\mathscr{S}_{R_{\alpha}}=0$ for every $90$ degrees of rotation, preserving the 4-fold rotational symmetry present in the underlying structure.
@@
This observation suggests that JS-SBM is less sensitive to the boundary characteristics and particle quantity within the finite cluster. Additionally, the fact that JS-SBM values range between 00 and $1$ facilitates direct comparison of SB across different clusters, making JS-SBM a more standardized metric. Consequently, the JS-SBM’s consistent and straightforward behavior makes it more suitable for comparative analysis of SB between distinct clusters compared to KL-SBM. Given these insights, the subsequent focus will exclusively be on the JS-SBM.
@@
By an inspection of the four-fold rotational symmetry, $C_{4}$ was not broken at all by the cubic and spherical cutouts but it was broken by the spheroidal and rectangular cutouts. For the sphere and cube, we find that $\mathscr{S}_{C_{4}}=1/3\sum_{i=1}^{3}\mathscr{S}_{C_{4}^{i}}=1/3\sum_{i=1}^{3}\mathscr{S}_{R_{i\pi/2}}=0$, and so our operator SBM correctly returns zero as expected. We find that for the spheroid, $\mathscr{S}^{JS}_{C_{4}}\approx 0.33$ and for the rectangular solid cutout, $\mathscr{S}^{JS}_{C_{4}}\approx 0.29$. Thus, by the measure $\mathscr{S}^{JS}_{C_{4}}$, we may argue that for this example, the spheroid cutout breaks the symmetry approximately the same as, but slightly more strongly than, the rectangular solid cutout.
5.1.2 The symmetry breaking from local perturbations
We now investigate the different ways that displaced atoms contribute to SB. As an illustrative example, we will use the same cubic cutout from the Nickel structure.
@@ 5.1.2 The symmetry breaking from local perturbations
We first test the four-fold rotational symmetry element $C_{4}$ that goes through the center of the cluster and is parallel to $c$, as previously discussed. Perturbations are introduced within the finite cluster by selecting a single atom and displacing it with a vector d. We then compute the JS-SBM, $\mathscr{S}^{JS}_{C_{4}}[\mu_{\textbf{d}}]$. The results are summarized in Fig. 3 plotted as a function of the magnitude of d, $\textbf{d}|$ d |.
Figure 3: The SBM $\mathscr{S}^{JS}_{C_{4}}[\mu_{d}]$ of a cubic cutout from the distorted Nickel structure. One of its face-centered atoms (darker blue atoms) is shifted parallel to the a-b plane (blue), deviated from the a-b plane by 45 degrees (red), or perpendicular to the a-b plane (yellow). The SBM is calculated for the four-fold rotational operation $C_{4}$, whose axis passes through the center of the finite cluster, with its direction along the positive c-axis. The curves with light color represent the SBM when the top or bottom face atom is displaced, while the darker curves represents the SBM when one of the other face atoms is displaced. The figure is plotted as a function of the length of the atom displacement $d$.
Displacing an atom lying on the rotation axis in a direction along the axis (top or bottom centering atom displaced along $c$) does not break symmetry and $\mathscr{S}^{JS}_{C_{4}}[\mu_{\textbf{d}}]=0$ for all $d$ (horizontal light yellow line in Fig. 3) as expected.
@@
Displacing these same atoms with a component perpendicular to the rotation axis does break symmetry. $\mathscr{S}^{JS}_{C_{4}}[\mu_{\textbf{d}}]$ increases smoothly as shown by the light blue and light red curves in Fig. 3. It saturates to a value of $0.072$ for all displacement directions. This value corresponds the atomic density of a single atom becoming disjoint, i.e. 1/14, since we have 14 atoms in the cluster. This value is reached when the atom moves far enough from its original position that it doesn’t overlap with itself or any other atoms on rotation.
@@
For displacements of atoms which are on the non-axial face centers, the SBM plateau is at $0.143$ ($\approx 2/14$). In this case the displaced atom is becoming disjoint with a neighboring atom and not just with itself.
We now compare in detail how the face-center atoms behave in three scenarios, for the displacement vectors being:
  1. 1.
    parallel to $c$ (shown in yellow).
  2. 2.
    within the $a$-$b$ plane (shown in blue) with two representative displacement directions $(1,0,0)$ and $(\sqrt{2}/2,\sqrt{2}/2,0)$).
  3. 3.
    at a $45^{\circ}$ angle to the $a$-$b$ plane (shown in red) with directions $(\sqrt{2}/2,0,\sqrt{2}/2)$ and $(1/2,1/2,\sqrt{2}/2)$.
@@
All the curves smoothly increase from 00 with increasing $d$ and eventually reach a plateau. We can gain insight into the behavior of the SBM by understanding this behavior. For the light blue curve, the atom on the rotation axis is displaced in the a-b plane by a distance $d$. After applying the $C_{4}$ symmetry operation to the dispaced atom, the displacement between the original and the rotated atom is $\sqrt{2}d$. The SBM will have its maximum value when the displaced but unrotated atomic density no longer overlaps with the displaced but rotated atomic density, which will occur at a $d=d^{*}/\sqrt{2}=0.93/\sqrt{2}.$
@@
Some of the atoms remain disjoint with increasing $d$ and $\mathscr{S}^{JS}_{C_{4}}[\mu_{d}]$ stays on the plateau, for example, the pale red and blue curves for the (1,0,0) displacement direction. However, in other cases, $\mathscr{S}^{JS}_{C_{4}}[\mu_{d}]$ becomes reduced with increasing $d$, an apparently paradoxical result that a larger distortion results in a lower SB. The reason for this is that the displacement of the atom is so large that it starts to overlap with another atom in the structure. For example, the displacement direction of (1,1,0) of the atom at the center of the top face when it reaches the corner of the cube. Logically and mathematically this makes sense, though this does not correspond to a real situation that would be encountered in practice as it is not possible for atomic probability distributions of pairs of atoms to overlap. However, plotting $\mathscr{S}^{JS}_{C_{4}}[\mu_{d}]$ over such and unphysically large range helps us to build intuition about the function. An analysis of SBM when displacing corner atoms is discussed in the Supplementary Materials.
@@
Figure 4: The SBM $\mathscr{S}^{JS}_{\sigma_{h}}[\mu_{d}]$ of a cubic cutout from the Nickel structure when one of its in-plane atoms (darker blue atoms) is shifted parallel to the plane (blue), deviated from the plane by 45 degrees (red), or perpendicular to the plane (yellow). The SBM is calculated for the reflection operation $\sigma_{h}$, whose mirror plane passes through the center of the finite cluster, with its normal vector along the positive $z$-axis. The figure is plotted as a function of the length of the atom displacement $d$.
@@
For an atom lying in the reflection plane atom, when the displacement is parallel to the mirror plane, $\mathscr{S}_{\sigma_{h}}[\mu_{d}]=0$ regardless of the displacement magnitude $d$, which is consistent with the blue curve in Fig. 4.
In contrast, when displacing an in-plane atom orthogonally to the mirror plane (as depicted by the yellow curve), $\mathscr{S}^{JS}_{\sigma_{h}}[\mu_{\textbf{d}}]$ increases and rapidly converges, reaching a value of approximately $0.072$ when $d=0.60$. In this case the SBM reaches its disjoint plateau at a value of $D$.
When considering a displacement vector angled at 45 degrees from the plane (as illustrated by the red curve), in scenarios of small $d$ values, $\mathscr{S}^{JS}_{\sigma_{h}}[\mu_{\textbf{d}}]$ primarily depends on the $c$-component of the displacement vector. Thus the curve rises more slowly and reaches the plateau at $\cos\theta D$, where $\theta$ is the angle the displacement vector makes with the reflection plane normal.
@@
Figure 5: The SBM $\mathscr{S}^{JS}_{\sigma_{h}}[\mu_{d}]$ of a cubic cutout from the Nickel structure when one of its in-plane atoms (darker blue atoms) is shifted parallel to the plane (blue), deviated from the plane by 45 degrees (red), or perpendicular to the plane (yellow). The SBM is calculated with respect to the reflection operation $\sigma_{h}$, whose mirror plane passes through the center of the finite cluster, with its normal vector along the positive $c$-axis. The figure is plotted as a function of the length of the atom displacement $d$ and focuses on smaller displacement magnitudes. The left-side red curve is derived by compressing the $d$-axis of the right-hand side red curve by a factor of $\sqrt{2}$.
As before, some atoms exhibit the anomalous reduction in SBM at large displacements. For instance, when an in-plane atom is displaced towards and overlaps one of its closest corner atoms, a local minimum is observed at $d=2.49$, characterized by $\mathscr{S}^{JS}{\sigma_{h}}[\mu_{\textbf{d}}]=0.018$. Further discussion can be found in Supplementary Materials. Moreover, an analysis of the sample size utilized for calculating the SBM is also available in the Supplementary Materials for further reference, as well as some computational details about computing the measure.
5.2 The violation of symmetry elements of local distorted perovskites from rotations of rigid units
@@
The collective rotations are modeled using algebraic expressions that link displacements of atoms so as to preserve the rigid linked octahedral rotations. We have employed similar algebraic expressions previously for data simulation [21]. Specifically we consider CaTiO3 and use crystallographically reasonable $U_{iso}$ values of $0.0052$2, $0.0027$2, and $0.0104$2 for Ca, Ti and O, respectively. The lattice parameter of the undistorted cubic perovskite is $3.91$. Illustrations of this in-phase tilt pattern, as viewed down each tilt axis, are shown in Figure 6.
@@
As this distortion process occurs, the space group changes from $Pm\bar{3}m$ (No. 221) to $P4/mbm$ (No. 127). Several symmetry operations that the tilted perovskite $\mu_{\phi}$ breaks and preserves are tested, and the evolution of the SBM $\mathscr{S}^{JS}_{\mathscr{O}}[\mu_{\phi}]$ is analyzed as the tilted angle around the $c$ axis, $\phi$, increases from 00 to $20$ degrees.
@@
Since the octahedral tilts do not disrupt the inversion symmetry, $\mathscr{S}^{JS}_{i}[\mu_{\phi}]$ maintains a value of zero for all $\phi$ as expected, illustrated by the constant line in Figure 7(c).
@@
6 Availability of code
The code for calculating the SBM is available as a Python-based open package in the following repository: https://github.com/diffpy/diffpy.sbm. The package is user-friendly, and detailed instructions for calculating the SBM for any finite cluster are provided in the repository’s README. Users can easily define their own finite clusters, or if working with crystals, the package supports importing unit cell data from a CIF file. Once the finite cluster is defined, you can apply distortions and calculate the SBM for various symmetry operations, such as rotation, reflection, inversion, and more.
Additionally, the code used to produce the results shown in this paper is publicly accessible at the GitHub repository: https://github.com/lanikaling/SymmetryBreakingMeasure, allowing readers to replicate the results or use it as a reference for their own work.
7 Conclusion

### v2

@@ V.1.1 The symmetry breaking from the boundary of the finite cluster
  • Spheroidal Cutout: Centered on a central atom, oriented with the unique long axis in the crystallographic $ab$-plane, extending two unit cells along the in-plane major axis, one unit cell along the orthogonal in-plane minor axis, and one unit cell along the $c$-axis, containing all atoms within this volume.
@@
In this test, the lattice parameter for nickel was chosen to be $3.52$ Å. The atomic displacement parameters (ADP) for each nickel atom were set to $U_{\text{iso}}$ =0.013=0.013 Å2. For each nickel atom, the electron count is 28, which aligns with its atomic number, $Z$. Additionally, the occupancy, $o$, is set to one for each site. For the evaluation of SBM, both $\mathscr{S}^{KL}_{R_{\alpha}}[\mu]$ and $\mathscr{S}^{JS}_{R_{\alpha}}[\mu]$ were estimated using Monte Carlo random sampling. The sample sizes for these estimations were determined based on a 95% confidence interval, with a bilateral error tolerance set at $0.025$ for the KL-SBM and $0.0025$ for the JS-SBM.
@@
We first consider the point-symmetry-preserving cutouts in Fig. 2 (a and c). By definition, both KL-SBM and JS-SBM $\mathscr{S}_{R_{\alpha}}[\mu]=0$ when $\alpha=0$ as the object is compared with itself. As we rotate the angle away from zero, $\mathscr{S}_{R_{\alpha}}[\mu]$ increases. Because the object has a four-fold rotational symmetry, $C_{4}$, for rotations about this axis, we expect both SBM $\mathscr{S}_{R_{\alpha}}[\mu]=\mathscr{S}_{C_{4}^{m}}[\mu]=0$ when $\alpha=m\pi/2$ rad for $m\in\{1,2,3\}$, which results in $\mathscr{S}_{R_{\alpha}}$ being a function with 4-fold periodicity in $\alpha$, as is seen. The fact that these cutouts do not break the point-symmetry is evident in the fact that $\mathscr{S}_{R_{\alpha}}=0$ for every $90$ degrees of rotation, preserving the 4-fold rotational symmetry present in the underlying structure.
@@
This observation suggests that JS-SBM is less sensitive to the boundary characteristics and particle quantity within the finite cluster. Additionally, the fact that JS-SBM values range between 0 and $1$ facilitates direct comparison of SB across different clusters, making JS-SBM a more standardized metric. Consequently, the JS-SBM’s consistent and straightforward behavior makes it more suitable for comparative analysis of SB between distinct clusters compared to KL-SBM. Given these insights, the subsequent focus will exclusively be on the JS-SBM.
@@
By an inspection of the four-fold rotational symmetry, $C_{4}$ was not broken at all by the cubic and spherical cutouts but it was broken by the spheroidal and rectangular cutouts. For the sphere and cube, we find that
$\mathscr{S}_{C_{4}}=\frac{1}{3}\sum_{m=1}^{3}\mathscr{S}_{C_{4}^{m}}=\frac{1}{3}\sum_{m=1}^{3}\mathscr{S}_{R_{m\pi/2}}=0,$ (18)
so our operator SBM correctly returns zero as expected. We find that for the spheroid, $\mathscr{S}^{JS}_{C_{4}}\approx 0.33$ and for the rectangular solid cutout, $\mathscr{S}^{JS}_{C_{4}}\approx 0.29$. Thus, by the measure $\mathscr{S}^{JS}_{C_{4}}$, we may argue that for this example, the spheroid cutout breaks the symmetry approximately the same as, but slightly more strongly than, the rectangular solid cutout.
V.1.2 The symmetry breaking from local perturbations
We now investigate the different ways that displaced atoms contribute to SB. As an illustrative example, we will use the same cubic cutout from the Nickel structure. Unless otherwise noted, we use the same isotropic ADP as above, $U_{\text{iso}}$ =0.013=0.013 Å2, for all nickel atoms.
@@ V.1.2 The symmetry breaking from local perturbations
We first test the four-fold rotational symmetry element $C_{4}$ that goes through the center of the cluster and is parallel to $c$, as previously discussed. Perturbations are introduced within the finite cluster by selecting a single atom and displacing it with a vector d. We then compute the JS-SBM, $\mathscr{S}^{JS}_{C_{4}}[\mu_{\textbf{d}}]$. The results are summarized in Fig. 3 plotted as a function of the magnitude of d, $d=|\textbf{d}|$. To guide the interpretation of Fig. 3, we consider three displacement-vector scenarios (in both panels):
  • within the $a$-$b$ plane (shown in blue) with two representative directions $(1,0,0)$ and $(\sqrt{2}/2,\sqrt{2}/2,0)$,
  • at a $45^{\circ}$ angle to the $a$-$b$ plane (shown in red) with two representative directions $(\sqrt{2}/2,0,\sqrt{2}/2)$ and $(1/2,1/2,\sqrt{2}/2)$,
  • parallel to $c$ (shown in yellow).
Panel a (left) shows displacements of a top/bottom face atom on the $C_{4}$ axis, whereas Panel b (right) shows displacements of a side-face atom off the axis.
Figure 3: JS–SBM $\mathscr{S}^{JS}_{C_{4}}[\mu_{\textbf{d}}]$ vs. displacement magnitude $d=|\textbf{d}|$. Panel a (left): displacing a top/bottom face atom on the $C_{4}$ axis. Colors encode displacement direction: blue (along $a$-$b$ plane), red ($45^{\circ}$), yellow (along $c$). Panel b (right): displacing a side-face atom off-axis (same color coding). The $C_{4}$ axis is fixed through the cluster center.
@@
Displacing an atom lying on the rotation axis in a direction along the axis (top or bottom centering atom displaced along $c$) does not break symmetry and $\mathscr{S}^{JS}_{C_{4}}[\mu_{\textbf{d}}]=0$ for all $d$ (yellow curve in Panel a of Fig. 3) as expected.
@@
Displacing these same atoms with a component perpendicular to the rotation axis does break symmetry. $\mathscr{S}^{JS}_{C_{4}}[\mu_{\textbf{d}}]$ increases smoothly as shown by the blue and red curves in Panel a of Fig. 3. It saturates to a value of $0.072$ for all displacement directions. This value corresponds to the atomic density of a single atom becoming effectively unmatched under the $C_{4}$ operation (i.e., $\approx 1/14$, since we have 14 atoms in the cluster). This plateau is reached when the atom moves far enough from its original position that the overlap under $C_{4}$ rotation is negligible.
@@
For displacements of atoms which are on the non-axial face centers, the SBM plateau is at $0.143$ ($\approx 2/14$). In this case the displaced atom becomes effectively unmatched with respect to both its own rotated copy and a neighboring site, rather than just itself (see Panel b).
@@
All the curves smoothly increase from 0 with increasing $d$ and eventually reach a plateau in both panel a and b. To gain insight into the behavior of the SBM, consider the blue curve in Panel a: the atom on the rotation axis is displaced in the $a$-$b$ plane by a distance $d$. After applying the $C_{4}$ symmetry operation to the displaced atom, the separation between the original and the rotated atom is $\sqrt{2}d$. The SBM approaches its plateau once the displaced but unrotated atomic density has a negligible overlap with its $C_{4}$-rotated counterpart. For Nickel’s isotropic ADP $U_{\mathrm{iso}}=0.013$ Å2, the single-atom translation case reaches near-saturation (JS–SBM $\gtrsim 0.95$) at a displacement of about $0.93$ Å in our simulation. In the present $C_{4}$ geometry the separation between those two positions is $\sqrt{2}\,d$, so the corresponding near-saturation in Panel a (blue curve) occurs at $d\approx 0.93/\sqrt{2}=0.66$ Å. This numerical value $0.93$ Å is specific to $U_{\mathrm{iso}}=0.013$ Å2. Larger $U_{\mathrm{iso}}$ requires a larger displacement to reach the same near-saturation level because overlap decays more slowly.
@@
Some of the atoms remain effectively non-overlapping under $C_{4}$ with increasing $d$ and $\mathscr{S}^{JS}_{C_{4}}[\mu_{\textbf{d}}]$ stays on the plateau; for example, the red and blue curves in Panel a for the $(1,0,0)$ displacement direction. However, in other cases, $\mathscr{S}^{JS}_{C_{4}}[\mu_{\textbf{d}}]$ becomes reduced with increasing $d$, an apparently paradoxical result that a larger distortion results in a lower SB. The reason for this is that the displacement of the atom is so large that it starts to regain appreciable overlap with another atom in the structure. For example, the displacement direction of $(1,1,0)$ of the atom at the center of the top face when it approaches the corner of the cube. Logically and mathematically this makes sense, though this does not correspond to a real situation that would be encountered in practice. However, plotting $\mathscr{S}^{JS}_{C_{4}}[\mu_{\textbf{d}}]$ over such an unphysically large range helps us to build intuition about the function.
@@
Figure 4: The SBM $\mathscr{S}^{JS}_{\sigma_{h}}[\mu_{\textbf{d}}]$ of a cubic cutout from the Nickel structure when one of its in-plane atoms (darker blue atoms) is shifted parallel to the plane (blue), deviated from the plane by 45 degrees (red), or perpendicular to the plane (yellow). The SBM is calculated for the reflection operation $\sigma_{h}$, whose mirror plane passes through the center of the finite cluster, with its normal vector along the positive $z$-axis. The figure is plotted as a function of the length of the atom displacement $d$.
@@
For an atom lying in the reflection plane atom, when the displacement is parallel to the mirror plane, $\mathscr{S}_{\sigma_{h}}[\mu_{\textbf{d}}]=0$ regardless of the displacement magnitude $d$, which is consistent with the blue curve in Fig. 4.
In contrast, when displacing an in-plane atom orthogonally to the mirror plane (as depicted by the yellow curve), $\mathscr{S}^{JS}_{\sigma_{h}}[\mu_{\textbf{d}}]$ increases and rapidly converges, reaching a value of approximately $0.072$ when $d=0.60$ Å. At this displacement, the atomic density and its reflection are effectively disjoint.
When considering a displacement vector angled at $45$ degrees from the plane (as illustrated by the red curve), the $\mathscr{S}^{JS}_{\sigma_{h}}[\mu_{\textbf{d}}]$ primarily depends on the component of the displacement vector normal to the plane, $d_{\perp}=d\cos\theta$, where $\theta$ is the angle between the displacement vector and the normal. Consequently, the curve rises more slowly compared to the orthogonal case.
@@
Figure 5: The SBM $\mathscr{S}^{JS}_{\sigma_{h}}[\mu_{\textbf{d}}]$ of a cubic cutout from the Nickel structure calculated for the reflection operation $\sigma_{h}$ (mirror plane normal to positive $c$-axis). The measure is plotted against displacement magnitude $d$ for shifts parallel to the plane (blue), perpendicular to the plane (yellow), and at 45 degrees (red). The yellow curve (left) represents the perpendicular displacement. It coincides with the trajectory obtained by scaling the $d$-axis of the 45-degree red curve (right) by a factor of $\cos(45^{\circ})=1/\sqrt{2}$, confirming that the scaled angled displacement is equivalent to the perpendicular displacement.
As before, some atoms exhibit the anomalous reduction in SBM at large displacements. For instance, when an in-plane atom is displaced towards and overlaps one of its closest corner atoms, depending on the displacement vector, a local minimum is observed at $d=2.49$ Å, characterized by $\mathscr{S}^{JS}_{\sigma_{h}}[\mu_{\textbf{d}}]=0.018$.
Monte Carlo sample size analysis
We make a note here to highlight certain numerical intricacies encountered during our tests. As discussed, we employ a Monte Carlo simulation to estimate the SBM. We found that the sample size required to obtain an estimate of the SBM at a certain level of accuracy increases as $(T_{\alpha})_{\#}\mu$ deviates further from $\mu$. This implies that as atomic displacements magnify, Monte Carlo calculations demand more extensive sample sizes to yield estimates of SBM with consistent precision. In Fig. 6, we plot the sample sizes that were used to estimate $\mathscr{S}^{JS}_{\sigma_{h}}[\mu]$. These sizes were determined by considering a $95\%$ confidence interval and a bilateral error tolerance of $0.0025$. The procedure for computing the sample size is delineated in Algorithm 2. All the simulation tests addressed in this section employ a sample size determined through this methodology.
Figure 6: The recommended sample size for Monte Carlo simulations of $\mathscr{S}^{JS}_{\sigma_{h}}[\mu]$. The sample size for (a) finite clusters with displacements of in-plane atoms and (b) for displacements of out-of-plane atoms. The colors of the curves indicate the sample size for displacements parallel to the plane (blue), 45 degrees to the plane (red), and perpendicular to the plane (yellow). These were the sample sizes needed to give $\mathscr{S}^{JS}_{\sigma_{h}}[\mu]$ with a $95\%$ confidence interval and an error tolerance of .0025.0025 on each side.
V.2 The violation of symmetry elements of local distorted perovskites from rotations of rigid units
@@
The collective rotations are modeled using algebraic expressions that link displacements of atoms so as to preserve the rigid linked octahedral rotations. Note that to maintain corner-connectivity, this distortion couples the tilt angle to the unit cell dimensions; consequently, the lattice parameters change, resulting in a non-cubic supercell and inducing positional shifts for all atoms in the finite cluster relative to the reference frame. We have employed similar algebraic expressions previously for data simulation [25].
Specifically we consider CaTiO3 and use crystallographically reasonable $U_{iso}$ values of $0.0052$ Å2, $0.0027$ Å2, and $0.0104$ Å2 for Ca, Ti and O, respectively. The lattice parameter of the undistorted cubic perovskite is $3.91$ Å. Illustrations of this in-phase tilt pattern, as viewed down each tilt axis, are shown in Figure 7.
@@
As this distortion process occurs, the space group changes from $Pm\bar{3}m$ (No. 221) to $P4/mbm$ (No. 127). Several symmetry operations that the tilted perovskite $\mu_{\phi}$ breaks and preserves are tested, and the evolution of the SBM $\mathscr{S}^{JS}_{\mathscr{O}}[\mu_{\phi}]$ is analyzed as the tilted angle around the $c$ axis, $\phi$, increases from 0 to $20$ degrees.
@@
We finally consider the inversion operator, $i$, defined with its fixed point at the central Ti cation. Since the octahedral tilts do not disrupt this inversion symmetry, $\mathscr{S}^{JS}_{i}[\mu_{\phi}]$ maintains a value of zero for all $\phi$ as expected, illustrated by the constant line in Figure 8(c).
@@
V.3 Extensions and potential applications
In the numerical tests presented above, we constructed the density $\mu$ using the Independent Atom Model (IAM), where each atom is represented by a Gaussian distribution derived from the average crystallographic ADPs ($U_{\text{iso}}$). This approach effectively captures the symmetry breaking of the crystallographically averaged structure. However, it does not account for correlated disorder.
Two-body correlations are experimentally accessible through the atomic pair distribution function (PDF) obtained from total-scattering experiments [5]. Accounting for such correlations could lead to different effective symmetry-breaking behavior, in which the symmetry breaking measure (SBM) returns different values depending on the range of interatomic distances over which it is evaluated. This situation arises, for example, in materials containing statistically disordered, fluctuating, locally symmetry-broken domains [3], where the local, low-$r$ symmetry may be lower than the domain-averaged symmetry observed at higher $r$.
The SBM is also affected by correlated atomic motions at short distances [15, 4]. Due to such correlations (e.g., rigid-bond behavior), the distribution of instantaneous interatomic distances is often significantly narrower than that predicted by the convolution of uncorrelated thermal ellipsoids. This leads to smaller effective static displacements and, consequently, to larger changes in the SBM at small interatomic distances.
The JS-SBM framework presented here is not limited to the analytic IAM representation. Since the Jensen-Shannon divergence is defined for general probability densities, the method can be extended to incorporate correlated motion and dynamic effects. While the PDF is a two-point correlation function rather than a probability density, it is a continuous function in $\mathbb{R}^{3}$, and a valid probability distribution can be constructed through appropriate normalization. For example, we propose
$P_{g}(r)=\frac{4\pi r^{2}\rho(r)}{\int_{0}^{r_{\max}}4\pi r^{2}\rho(r)\,\mathrm{d}r},$ (19)
where $r_{\max}$ is the maximum distance considered, and $\rho(r)$ is the atomic-pair density defined as the number of atoms found in a spherical shell of thickness $\mathrm{d}r$ at a distance $r$ from a reference atom, averaged over all atoms in the sample [8, 5]. This quantity can be computed directly from structural models and, notably, from the experimentally accessible $G(r)$ function via
$\rho(r)=\frac{G(r)}{4\pi r}-\gamma_{0}(r)\rho_{0}.$ (20)
Here $\rho_{0}$ is the average atomic density, and $\gamma_{0}(r)$ is the characteristic function of the sample. It takes a value of unity for an infinite crystal and, for finite particles, represents the orientationally averaged autocorrelation of the particle shape function, giving the probability that two points separated by a distance $r$ both lie within the particle.
Although we have not carried out a detailed analysis of the SBM applied to this distribution, such a calculation would be straightforward to perform, for example, on ensembles of instantaneous configurations generated using Molecular Dynamics (MD) or Reverse Monte Carlo (RMC) simulations. In this context, the SBM would quantify instantaneous symmetry breaking, capturing local distortions that may average out in time-averaged structures but remain visible in PDF measurements. Moreover, since $G(r)$ is directly accessible experimentally, this PDF-based SBM could be computed without first solving the structure, for example in a temperature-dependent study across a structural phase transition.
The ability to define SBMs for different probability distributions also suggests that, more generally, a family of symmetry-breaking measures may be constructed for material systems, potentially revealing insights that are not accessible through a single representation. For example, Fang et al. [6] introduced a continuous symmetry measure that quantifies changes in a system’s Hamiltonian, rather than the downstream effects on the IAM or the PDF.
This flexibility suggests that the SBM could serve as a valuable tool within modeling frameworks such as RMCProfile [28] or PDFgui [7, 33]. By incorporating the SBM as a continuous order parameter or as a regularization term in the loss function, researchers could quantitatively monitor specific symmetry violations during refinement, enabling a more controlled exploration of disorder and local symmetry breaking.
VI Availability of code
The software code used to produce the results shown in this paper, including the implementation of the SBM algorithms, is publicly accessible at the GitHub repository: https://github.com/lanikaling/SymmetryBreakingMeasure. This repository contains detailed instructions for calculating the SBM for general finite clusters. Users can manually define their own finite clusters or import unit cell data directly from a CIF file. The code supports applying distortions and calculating the SBM for various symmetry operations, such as rotation, reflection, and inversion. Furthermore, we plan to eventually integrate this functionality into the DiffPy organization as a dedicated package (https://github.com/diffpy/diffpy.sbm) to provide a standardized open-source distribution in the future.
VII Conclusion